# Image Inpainting based on Visual-Neural-Inspired Specific Object-of-Interest Imaging Technology

### v2

@@ 3.1 Specific Object-of-Interest Imaging
Assume the dictionary $\Psi$ satisfies the Restricted Isometry Property (RIP), i.e. $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ (separability condition). The underlying mechanism is that the filter $\phi$ implements an (approximate) orthogonal complement in feature space, $\phi=I-P_{\mathrm{bg}}$, thereby suppressing non-target components (i.e., those with $i\neq 0$).
A closed-form estimator for $\phi$ can be obtained by minimizing a least-squares criterion and expressed via the Moore–Penrose pseudoinverse. Defining $k^{(h)}$ as the vectorized scene for sample $h$ and $\widehat{k}_{0}^{(h)}$ as its corresponding target component, we may write
@@
where denotes the Moore–Penrose pseudoinverse, and $k^{(h)}\in\mathbb{R}^{D\times 1}$ is the vectorized scene representation for sample $h$. We further constrain $\phi=\mathbf{I}-\mathbf{P}_{\mathrm{bg}}$ as an (approximate) orthogonal projection onto the target subspace.
Equivalently, $\phi$ may be viewed as the best approximation to the ideal projection $I-P_{\mathrm{bg}}$ in a suitable Hilbert space $\mathcal{H}$:
@@
Here $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$; $\|\cdot\|_{F}$ denotes the Frobenius norm for matrices, $\|A\|_{F}=\sqrt{\sum_{i}\sum_{j}|a_{ij}|^{2}}$.
Let $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ denote learned query and key feature vectors at spatial locations $(x,y)$ and $(i,j)$, respectively, and let $d$ be the feature dimensionality used for softmax scaling. Under bio-inspired attentional modulation we compute attention weights by
@@
where the denominator normalizes over all spatial locations $(i,j)$. When $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ are highly similar, the corresponding attention weight $\alpha_{xy}$ approaches one, consistent with Feature Integration Theory: biased competition in V4 enhances target features via attentional gain.
Require that the Kullback–Leibler divergence between the target and background distributions satisfies
@@
where $D_{\mathrm{KL}}(P\|Q)=\sum_{i}P_{i}\log\frac{P_{i}}{Q_{i}}$ denotes the Kullback–Leibler divergence.
By large-deviation principles, when the distributional discrepancy between the target and background dominates inter-object interference, empirical distributions concentrate on distributions minimizing relative entropy; consequently, the estimator tends to concentrate on the target distribution, which implies reduction of the estimated target entropy under suitable conditions. Hence, under these assumptions there exists a unique solution $\widehat{k}_{0}$ such that
@@
where $H(X)=-\sum_{x\in\mathcal{X}}p(x)\log p(x)$ denotes the Shannon entropy.
Under Assumptions A1–A3 (sparsity of object representations, the RIP on dictionary $\Psi$, and bounded measurement noise), there exists a linear operator $\phi$ (learnable by FECNet) that approximates the ideal target projector $I-P_{\mathrm{bg}}$. By compressed-sensing theory, if $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ then the learned projector $\phi$ (approximating $I-P_{\mathrm{bg}}$) admits the error bound
@@
where $C_{1}$ and $C_{2}$ depend on the RIP constants and the dictionary condition number; $\sigma_{s}(k)$ denotes the error of the best $s$-term approximation of $k$; and $\epsilon$ denotes an upper bound on the measurement noise.
Together, these results provide a rigorous foundation for the proposed object-centric imaging approach. By casting target separation as a formal projection problem under sparsity constraints and employing information-theoretic concentration arguments to characterize convergence, we show that under the stated assumptions the Specific Object-of-Interest Imaging system provably isolates the target and converges to an optimal reconstruction that up to the approximation and noise terms given above.
@@ 3.2 Image Inpainting System
To enhance structural coherence and reconstruction accuracy for occluded images, we propose a two-stage framework grounded in specific-object-of-interest imaging principles (Fig. 3). Departing from traditional pixel-level completion in RGB space, our architecture adopts an object-first perception inspired by biological vision, decomposing the task into structure-prioritized progressive refinements. Emphasizing semantically coherent reconstruction, we decompose the imaging pipeline into two stages: (1) object perception $\rightarrow$ (2) object inpainting. The core innovation is SIOI, which compels the model to focus on the target object while suppressing irrelevant background interference. This reduces computational overhead by concentrating structural modeling on critical regions (see Fig. 4).
@@
Stage I: Specific Object-of-Interest Imaging Biological vision begins with retinal photoreceptors converting incoming light into neural signals. Subsequent processing in the lateral geniculate nucleus (LGN) integrates spatial information and suppresses high-frequency noise via center–surround antagonism. Feature extraction proceeds in primary visual cortex (V1), where simple cells respond to oriented edges and spatial-frequency textures [46]. Object prioritization amplifies target representations through bottom-up attentional biasing [66], suppressing background interference—functionally analogous to selective attention in the ”cocktail-party” effect.
Our SIOI module computationally embodies this hierarchy. We process corrupted inputs with modified U-Net encoders to extract multi-scale features, mimicking V1 orientation selectivity. Embedding Fast Fourier Convolution (FFC) modules in multiple encoding and decoding processes captures long-range dependencies through spectral transformation, simulating attentional modulation in V4 that amplifies target signals. The spectral processing can be written as
@@
where $\mathrm{FFT2}$ and $\mathrm{IFFT2}$ denote the 2-D Fourier transform and its inverse, respectively, and $\mathrm{Conv}_{\mathrm{freq}}$ is a learnable convolutional operator applied in the frequency domain.(Fig. 5)
The output of the SIOI module is a preliminary, yet structurally coherent, object map contain local defects. As shown in Fig. 4. Biologically, this parallels V1 edge detection and V4-mediated object-based attention, in which attentional gain enhances target representations while suppressing background activity [74]. Our FFC modules enhance target-relevant spectral components while attenuating background spectra. Functionally, FFC operations correspond to V1-like orientation detection, whereas the learned attention weights emulate pulvinar–thalamic modulation of cortical processing. Consequently, Stage I produces structurally coherent object maps despite occlusions, providing essential priors for subsequent texture synthesis and refinement.This spectral processing overcomes the limitations of standard convolutions for modeling long-range dependencies and significantly improves structural consistency.
Together, these mechanisms reduce information redundancy. The SIOI module serves as a generic, front-end ”visual processor” that transforms a complex, occluded scene into a purified, object-centric representation. It provides a powerful intermediate representation that is the cornerstone of our framework, explicitly containing the structural and semantic priors necessary for high-fidelity inpainting.
Stage II: Object Inpainting (Coarse inpainting + Fine inpainting) To validate the efficacy of the SIOI module independently of advanced inpainting architectures, we deliberately employ a straightforward inpainting network in Stage II. Unlike conventional image inpainting methods, our model eliminates the need for additional manually annotated corruption masks and is capable of automatically identifying and inpainting the corrupted regions of the target. It is critical to emphasize that the innovation of this work is not this network itself, but rather the SIOI module. The purpose of Stage II is solely to demonstrate that the priors generated by SIOI are so potent that they enable even a simple network to achieve high-quality reconstruction. This stage is loosely analogous to the role of higher visual areas in integrating semantic information for detailed perception. Our Structural Recovery Module (SRM) takes the SIOI output and performs contour interpolation and shape completion using multi-scale dilated convolutions and feature-pyramid fusion:
@@
An FPN-like fusion integrates encoder–decoder hierarchical features, and subpixel-convolution upsampling is used to suppress checkerboard artifacts. The output is a complete structural feature tensor $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$.
@@
Functionally, both biological and computational systems achieve structural coherence: biological vision generates shape priors via predictive-coding [50], while the SRM fuses multi-scale context through MSDC and feature-pyramid fusion. This design helps to avoid texture blurring and semantic mismatches. Technically, MSDC operations correspond to V2–V3 contour integration, while feature-pyramid fusion approximates cortical feedback modulation. SRM enhances structural integrity in partially occluded object maps from Stage I, enabling a coarse-to-fine imaging strategy.
@@
where $\mathcal{R}(\cdot)$ denotes SRM processing and $I_{\mathrm{coarse}}$ denotes the coarsely restored map. Subsequently, the Global Detail Recovery Module (GDRM) utilizes a symmetric encoder-decoder with skip connections to synthesize fine-grained textures onto the recovered structure, guided by the robust priors from Stage I. Our computational counterpart, the Global Detail Recovery Module, implements a symmetric encoder–decoder architecture that consumes the 64-channel structural features $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$ produced by the SRM. This simple design choice underscores our core claim: The SIOI module is a versatile, plug-and-play component that can elevate the performance of any existing inpainting pipeline by providing superior object-specific guidance.
@@ 3.3 Loss function
To achieve high-fidelity reconstruction of defective object images, we design a multi-branch fused loss that jointly promotes structural coherence, fine-detail reconstruction, and global visual-style consistency. The composite loss comprises three components: a weighted reconstruction loss [48, 75], a perceptual loss [76], and a style loss [77]. Integrated with principled weighting and multi-stage optimization strategies, the loss ensures that synthesized images exhibit improved semantic structure and textural fidelity.
Weighted Reconstruction Loss. We introduce a spatially weighted reconstruction loss based on target maps. An explicit weight map $w(x,y)$ differentially weights per-pixel reconstruction errors in the $L_{1}$$L_{2}$ terms:
@@
where $\hat{I}(x,y)$ denotes the predicted intensity at pixel $(x,y)$, $I_{\mathrm{gt}}(x,y)$ the corresponding ground truth, $I_{i}^{\mathrm{pred}}$ and $I_{i}^{\mathrm{gt}}$ the RGB vectors at pixel $i$, $w_{i}$ the pixel-level weight, and $\alpha$ a scalar weighting factor (empirically set to 4). Higher weights are assigned to semantic target regions and lower weights to background regions; this directs the network to prioritize detail reconstruction in structurally critical areas and effectively mitigates edge blurring and structural discontinuities. This mechanism intensifies supervision in target regions while reducing gradient interference from irrelevant areas.
Perceptual Loss. To compensate for pixel-level losses’ inability to capture high-level semantics, we incorporate a perceptual loss computed from feature activations of a pretrained VGG-16 network:
@@
where $\phi_{l}(\cdot)$ denotes the feature maps extracted from the $l$-th VGG-16 layer, and $C_{l},H_{l},W_{l}$ denote the channel count, height, and width of $\phi_{l}$, respectively. By comparing representations across selected VGG-16 layers, the perceptual loss captures high-level semantics (e.g., object shapes and textures) and enhances semantic coherence in generated objects.
Style Loss. The style loss quantifies texture distribution and stylistic characteristics using Gram matrices as in style-transfer methods [77]. For feature maps $\phi_{l}(x)$ at layer $l$ of VGG-16, the Gram matrix is
@@
where $\mathcal{L}$ denotes the set of selected layers and $\|\cdot\|_{F}$ the Frobenius norm. This statistical matching captures local texture statistics and global stylistic discrepancies, improving the naturalness of restored images and reducing pattern repetition and artifacts.
Composite Loss. The integrated loss combines the above terms. The weighted reconstruction term enhances edge clarity and structural continuity in semantic regions; the perceptual term reinforces high-level semantic consistency aligned with human perception; and the style term mitigates texture artifacts across diverse objects and scenes:
@@
These weights are empirically chosen to balance the differing numerical scales and competing objectives of the loss components during training.
@@ 4.1 Dataset
Four distinct datasets were employed. A custom object-centric teapot dataset (2,000 training, and 200 test samples) constructed in our laboratory featuring $512\times 512$ scene images of multi-view teapot objects against complex backgrounds with artificial items and indoor environments, each paired with corresponding $512\times 512$ ground-truth masks that underwent multi-stage manual calibration to preserve geometric structures and textural details; alongside three COCO2017-derived subsets—“elephant” (2,143 training images, 89 validation as test), “giraffe” (2,546 training, 101 validation as test), and “zebra” (1,916 training, 85 validation as test)—all uniformly resized to $512\times 512$ resolution to ensure comparability. A consistent target-aware occlusion protocol was implemented across all datasets, automatically generating masks from target maps to restrict occlusions exclusively to foreground areas, with coverage constrained to $\sim$40% of object regions during both training and inference phases to preserve contextual information while necessitating complex structural reasoning.
@@ 4.2 Experimental Setup
Implementation Details. All experiments were conducted on an Ubuntu 20.04 system equipped with an Intel Xeon Platinum 8358P CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM). The framework was implemented in PyTorch 1.10.0 with CUDA 11.3 support. Training spanned 50 epochs with a batch size of 4, constrained by available computational resources. Optimization employed Adam with a learning rate of $2\times 10^{-4}$, resulting in approximately 0.5 hours of total training time on the specified hardware. We adopted a two-stage training protocol: First, the ObjectImagingModule was pretrained exclusively using $\mathcal{L}_{\mathrm{recon}}$ to generate semantically structured intermediate representations. This plug-and-play module was subsequently frozen and integrated across all four experimental configurations. Second, the Structural Refinement Module (SRM) and Geometric Detail Reconstruction Module (GDRM) were trained using $\mathcal{L}_{\mathrm{perc}}$ and $\mathcal{L}_{\mathrm{style}}$ to enhance semantic coherence and stylistic fidelity. This secondary optimization (implemented solely for preliminary validation in Experiment 1) maintained the ObjectImagingModule’s parameters fixed. Full implementation code is available at: https://github.com/WYH302/WuYonghao.git.
@@
Baseline. To rigorously evaluate the efficacy of our proposed method, we benchmark against a representative spectrum of state-of-the-art inpainting approaches selected through a multi-faceted lens. Our criteria prioritize comprehensive coverage of methodological evolution, spanning from classical exemplar-based techniques to cutting-edge generative models, while ensuring architectural diversity across diffusion, transformer, and GAN-based paradigms. The evaluation suite begins with foundational exemplar-based methods: Criminisi [18], representing classical non-local patch synthesis, establishes a performance baseline for traditional non-deep-learning approaches and highlights inherent limitations in complex structural recovery. Its optimized counterpart, PatchMatch [78], provides critical insights into computational efficiency trade-offs through accelerated nearest-neighbor matching. Transitioning to modern deep learning, AOT-GAN [79] is included for its attention-driven, long-range dependency modeling via dilated convolutions—directly addressing holistic scene coherence challenges central to our work. Edge-Connect [56] offers a strategic comparative lens as a two-stage edge-guided model, enabling explicit evaluation of structural prior utilization against our coarse-to-fine refinement paradigm. The hybrid methodology of MISF [80] tests multi-scale feature interaction between traditional and deep components against our biologically inspired framework. Finally, contemporary diffusion models LaMa [81] and RePaint [41] are incorporated as cutting-edge references, with LaMa validating spectral processing choices and RePaint contrasting stochastic refinement against our deterministic pipeline. All baselines underwent five independent runs with randomized initializations, utilizing official implementations and pre-trained models where available.
Evaluation metrics. We employed four widely used metrics to quantify inpainting performance: (1) SSIM (Structural Similarity Index Measure), which assesses structural similarity between predictions and ground truth (higher is better); (2) PSNR (Peak Signal-to-Noise Ratio), measuring pixel-wise reconstruction fidelity (higher is better); (3) MAE (Mean Absolute Error), which measures average absolute pixel deviation (lower is better); and (4) LPIPS [82], a learned perceptual metric based on deep features (lower is better).

### v3

@@ 3.1 Specific Object-of-Interest Imaging
Assuming the dictionary $\Psi$ satisfies the Restricted Isometry Property (RIP), i.e., $\mathrm{RIP}(\Psi)\leq\delta_{2s}$ (separability condition), the filter $\phi$ approximates an orthogonal complement in feature space ($\phi=I-P_{\mathrm{bg}}$), suppressing non-target components ($i\neq 0$).
A closed-form estimator for $\phi$ is derived by minimizing a least-squares criterion using the Moore-Penrose pseudoinverse:
@@
where denotes the Moore-Penrose pseudoinverse, and $k^{(h)}\in\mathbb{R}^{D\times 1}$ is the vectorized scene for sample $h$. We constrain $\phi=\mathbf{I}-\mathbf{P}_{\mathrm{bg}}$ as an approximate orthogonal projection onto the target subspace.
Equivalently, $\phi$ is the best approximation to $I-P_{\mathrm{bg}}$ in a Hilbert space $\mathcal{H}$:
@@
Here $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and norm $\|\cdot\|$; $\|\cdot\|_{F}$ is the Frobenius norm, where $\|A\|_{F}=\sqrt{\sum_{i}\sum_{j}|a_{ij}|^{2}}$.
Let $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ be learned query and key vectors at locations $(x,y)$ and $(i,j)$, respectively, with $d$ as the feature dimensionality for softmax scaling. Under bio-inspired attentional modulation, attention weights are computed as:
@@
where the denominator normalizes over all locations $(i,j)$. When $\mathbf{q}_{xy}$ and $\mathbf{k}_{ij}$ are similar, $\alpha_{xy}$ approaches 1, consistent with Feature Integration Theory: biased competition in V4 enhances targets via attentional gain.
We require the Kullback-Leibler (KL) divergence between target and background distributions to satisfy:
@@
where $D_{\mathrm{KL}}(P\|Q)$ is the KL divergence.
By large-deviation principles, if the target-background discrepancy dominates inter-object interference, empirical distributions concentrate on entropy-minimizing distributions. Thus, the estimator concentrates on the target distribution, reducing estimated target entropy under suitable conditions. Hence, under these assumptions, a unique solution $\widehat{k}_{0}$ exists such that:
@@
where $H(X)$ is the Shannon entropy.
Under Assumptions A1–A3 (object sparsity, RIP on $\Psi$, and bounded noise), a linear operator $\phi$ (learnable by FECNet) approximates the ideal projector $I-P_{\mathrm{bg}}$. By compressed-sensing theory, if $\mathrm{RIP}(\Psi)\leq\delta_{2s}$, the learned projector $\phi$ admits the error bound:
@@
where $C_{1}$ and $C_{2}$ depend on RIP constants and the dictionary condition number; $\sigma_{s}(k)$ is the best $s$-term approximation error; and $\epsilon$ bounds the measurement noise.
These results establish a rigorous foundation for our object-centric imaging approach. By formulating target separation as a sparsity-constrained projection problem and using information-theoretic arguments, we show that under our assumptions, the SIOI system provably isolates the target and converges to an optimal reconstruction, up to approximation and noise errors.
@@ 3.2 Image Inpainting System
We propose a two-stage framework based on specific-object-of-interest imaging to enhance structural coherence and reconstruction accuracy for occluded images (Fig. 3). Unlike traditional RGB-space pixel completion, our architecture adopts a biologically inspired object-first approach, decomposing the task into structure-prioritized refinements. We decompose the pipeline into: (1) object perception $\rightarrow$ (2) object inpainting, emphasizing semantic coherence. The core innovation, SIOI, directs model focus to the target while suppressing background interference. This reduces computation by concentrating on critical target regions (see Fig. 4).
@@
Stage I: Specific Object-of-Interest Imaging Biological vision starts with retinal phototransduction. Subsequent LGN processing integrates spatial information and suppresses high-frequency noise through center-surround antagonism. In primary visual cortex (V1), simple cells extract oriented edges and spatial-frequency textures [49]. Object prioritization amplifies targets via bottom-up attention [68], suppressing background interference—analogous to the ”cocktail-party” effect.
Our SIOI module computationally implements this hierarchy. Using modified Encoder-Decoder, we process corrupted inputs to extract multi-scale features, mimicking V1 orientation selectivity. FFC modules in encoding/decoding capture long-range dependencies via spectral transformation, simulating V4 attentional modulation that amplifies target signals. The spectral processing is:
@@
where $\mathrm{FFT2}$$\mathrm{IFFT2}$ are 2-D Fourier transform/inverse, and $\mathrm{Conv}_{\mathrm{freq}}$ is a learnable frequency-domain convolutional operator (Fig. 4).
The SIOI module outputs a preliminary but structurally coherent object map with local defects, as shown in Fig. 4. Biologically, this mimics V1 edge detection and V4-mediated attention, where attentional gain enhances targets and suppresses background activity [76]. Our FFC modules amplify target-relevant spectral components and attenuate background spectra. Functionally, FFC operations resemble V1 orientation detection, while attention weights emulate pulvinar-thalamic cortical modulation. Thus, Stage I generates structurally coherent object maps despite occlusions, providing essential priors for texture synthesis and refinement. This spectral processing improves upon standard convolutions by modeling long-range dependencies and enhancing structural consistency.
These mechanisms collectively reduce information redundancy. The SIOI module acts as a front-end visual processor, transforming complex, occluded scenes into purified, object-centric representations. It provides a key intermediate representation, containing structural and semantic priors essential for high-fidelity inpainting.
Stage II: Object Inpainting (Coarse + Fine) To validate SIOI module efficacy independently, we use a simple inpainting network in Stage II. Unlike conventional methods, our model requires no manual masks and automatically identifies and inpaints target corrupted regions. Critically, the innovation is the SIOI module, not this network. Stage II demonstrates that SIOI priors enable high-quality reconstruction even with a simple network. This stage analogizes higher visual areas integrating semantic information for detailed perception. Our Structural Recovery Module (SRM) processes SIOI output with multi-scale dilated convolutions and feature-pyramid fusion for contour interpolation and shape completion:
@@
FPN-like fusion integrates hierarchical features, and subpixel-convolution upsampling suppresses checkerboard artifacts. The output is a complete structural feature tensor $\mathcal{F}_{\mathrm{SRM}}\in\mathbb{R}^{64\times H\times W}$.
@@
Functionally, both systems achieve structural coherence: biological vision uses predictive-coding for shape priors [53], while SRM fuses multi-scale context via MSDC and feature-pyramid fusion. This design avoids texture blurring and semantic mismatches. Technically, MSDC corresponds to V2–V3 contour integration, and feature-pyramid fusion approximates cortical feedback. SRM enhances structural integrity in occluded maps from Stage I, enabling coarse-to-fine imaging.
@@
where $\mathcal{R}(\cdot)$ is SRM processing and $I_{\mathrm{coarse}}$ is the coarsely restored map. The Global Detail Recovery Module (GDRM) then uses a symmetric encoder-decoder with skip connections to synthesize fine textures, guided by Stage I priors. GDRM implements a symmetric encoder-decoder that consumes the 64-channel features $\mathcal{F}_{\mathrm{SRM}}$ from SRM. This design underscores that SIOI is a versatile, plug-and-play component that elevates any inpainting pipeline via superior object-specific guidance.
@@ 3.3 Loss function
To achieve high-fidelity reconstruction of defective object images, we propose a multi-branch composite loss function that simultaneously enhances structural coherence, fine-detail recovery, and global style consistency. It integrates three components: (1) a weighted reconstruction loss [51, 77], (2) a perceptual loss [78], and (3) a style loss [79]. With carefully designed weighting schemes and a multi-stage optimization strategy, this composite loss ensures semantically coherent structures and texturally faithful details in the synthesized images.
Weighted Reconstruction Loss. We first introduce a spatially weighted reconstruction loss guided by target-specific maps. An explicit weight map $w(x,y)$ differentially weights per-pixel reconstruction errors in the $L_{1}$$L_{2}$ terms:
@@
where $\hat{I}(x,y)$ denotes the predicted intensity at pixel $(x,y)$, $I_{\mathrm{gt}}(x,y)$ the corresponding ground truth, $I_{i}^{\mathrm{pred}}$ and $I_{i}^{\mathrm{gt}}$ the RGB vectors at pixel $i$, $w_{i}$ the pixel-level weight, and $\alpha$ a scalar weighting factor (empirically set to 4). This weighting scheme prioritizes detail reconstruction in semantically important regions while reducing emphasis on background areas, effectively mitigating edge blurring and structural discontinuities. This approach enhances supervision on target regions while minimizing interference from irrelevant background gradients.
Perceptual Loss. To address the limitation of pixel-level losses in capturing high-level semantics, we incorporate a perceptual loss derived from feature activations of a pretrained VGG-16 network:
@@
where $\phi_{l}(\cdot)$ denotes the feature maps extracted from the $l$-th VGG-16 layer, and $C_{l},H_{l},W_{l}$ denote the channel count, height, and width of $\phi_{l}$, respectively. By comparing deep feature representations across selected VGG-16 layers, the perceptual loss captures high-level semantic information (e.g., object shapes and textures) and improves semantic coherence in the generated outputs.
Style Loss. The style loss measures texture and style characteristics via Gram matrices, following style-transfer approaches [79]. For feature maps $\phi_{l}(x)$ at layer $l$ of VGG-16, the Gram matrix is
@@
where $\mathcal{L}$ denotes the set of selected layers and $\|\cdot\|_{F}$ the Frobenius norm. This statistical matching captures both local texture patterns and global style consistency, enhancing realism in restored images while reducing artifacts and repetitive patterns.
Composite Loss. Finally, we combine these components into a composite loss:
@@
The reconstruction term sharpens edges in key regions, the perceptual term ensures semantic alignment with human vision, and the style term reduces texture artifacts across varied content. These weights balance the varying scales and objectives of the loss components during training.
@@ 4.1 Dataset
We utilized four distinct datasets for our experiments. This included a custom, object-centric teapot dataset (2,000 training and 200 test samples), created in-house. It contains $512\times 512$‘ images of teapots from multiple viewpoints, set against complex backgrounds with artificial items and indoor settings. Each image has a corresponding $512\times 512$‘ ground-truth mask, meticulously calibrated through a multi-stage manual process to ensure accurate geometric and textural details. Additionally, three subsets were derived from COCO2017: “elephant” (2,143 training, 89 validation-test), “giraffe” (2,546 training, 101 validation-test), and “zebra” (1,916 training, 85 validation-test). All images were resized to $512\times 512$‘ to maintain consistency. A consistent occlusion protocol was applied across all datasets. Masks were automatically generated from target maps to ensure occlusions affected only the foreground objects, covering approximately $\sim$40%‘ of their area. This balance preserved sufficient context while requiring the model to perform non-trivial structural reasoning during both training and inference.
@@ 4.2 Experimental Setup
Implementation Details. All experiments were run on a Ubuntu 20.04 system with an Intel Xeon Platinum 8358P CPU and an NVIDIA GeForce RTX 4090 GPU (24 GB VRAM). We implemented our framework in PyTorch 1.10.0 using CUDA 11.3. Training lasted for 50 epochs with a batch size of 4, a constraint imposed by our computational resources. We used the Adam optimizer with a learning rate of $2\times 10^{-4}$‘. The total training time was approximately 0.5 hours on this hardware. Our training followed a two-stage protocol. First, we pre-trained the ObjectImagingModule using only the reconstruction loss $\mathcal{L}_{\mathrm{recon}}$‘ to produce structured intermediate representations. This plug-and-play module was then frozen and integrated into all subsequent experimental setups. Second, we trained the Structural Refinement Module (SRM) and Geometric Detail Reconstruction Module (GDRM) with the perceptual loss $\mathcal{L}_{\mathrm{perc}}$‘ and style loss $\mathcal{L}_{\mathrm{style}}$‘ to improve semantic coherence and stylistic fidelity. During this second stage, the ObjectImagingModule’s parameters remained fixed. This optimization was conducted primarily for preliminary validation in Experiment 1. Our full implementation code is available at: https://github.com/WYH302/WuYonghao.git.
@@
Baseline. To rigorously evaluate our method, we benchmark against a representative selection of state-of-the-art inpainting approaches. Our selection emphasizes methodological diversity, encompassing classical exemplar-based methods to modern generative approaches, including diffusion models, transformers, and GAN architectures. We include Criminisi [18] as a foundational exemplar-based method that uses non-local patch synthesis, providing a baseline for traditional approaches while revealing limitations in complex structural recovery. PatchMatch [80], an optimized variant, offers insights into computational efficiency through accelerated nearest-neighbor matching. From modern deep learning, AOT-GAN [81] is included for its attention-driven modeling of long-range dependencies using dilated convolutions, addressing holistic scene coherence. Edge-Connect [44] provides a two-stage edge-guided approach that allows comparison of structural prior usage with our coarse-to-fine framework. The hybrid MISF [82] tests multi-scale feature interactions between traditional and deep components, providing a contrast to our bio-inspired approach. Finally, we include contemporary diffusion models: LaMa [83] validates our spectral processing choices, while RePaint [41] provides a stochastic refinement contrast to our deterministic approach. All baseline methods were executed for five independent runs with random initializations, using official implementations and pre-trained models when available.
Evaluation metrics. We used four standard metrics: (1) SSIM measures structural similarity (higher values preferred), (2) PSNR quantifies pixel-level fidelity (higher values better), (3) MAE calculates average pixel deviation (lower values better), and (4) LPIPS [84] evaluates perceptual quality (lower values better).